A chemist dilutes a 32% saline (salt) solution with water (0% saline) in order to produce 150 liters of 20% saline solution. How much water and how much 32% solution does she need to use?
Thank you for taking the time to answer my question.

Answer by Paul(988)

Let x be the amount of salt used
Let y be the amount of water used
y=150-x (subsitution)
32x+0y=150(20)
Subsitute:
32x+0(150-x)=3000
32x=3000
x=93.75
y=150-93.75
y=56.25
Hence, a chemist uses 93.75 L of salt and 56.25L of water.
Paul.